### v2

# Lossless data compression at pragmatic rates

Abstract
The problem of variable-rate lossless data compression is considered, for codes with and without prefix constraints. Sharp bounds are derived for the best achievable compression rate of memoryless sources, when the excess-rate probability is required to be exponentially small in the blocklength. Accurate nonasymptotic expansions with explicit constants are obtained for the optimal rate, using tools from large deviations and Gaussian approximation. Examples are shown indicating that, in the small excess-rate-probability regime, the approximation to the fundamental limit of the compression rate suggested by these bounds is significantly more accurate than the approximations provided by either normal approximation or error exponents. The new bounds reinforce the crucial operational conclusion that, in applications where the blocklength is relatively short and where stringent guarantees are required on the rate, the best achievable rate is no longer close to the entropy. Rather, it is an appropriate, more pragmatic rate, determined via the inverse error exponent function and the blocklength.
00footnotetext: This work was supported in part by the EPSRC-funded INFORMED-AI project EP/Y028732/1.
Keywords — Data compression, memoryless source, error exponent, normal approximation, large deviations, pragmatic rates
1 Introduction
1.1 Variable-rate data compression
[…]
We revisit the problem of losslessly compressing the output of such a memoryless source. The problem formulation is simple and quite elementary, though its importance can hardly be overstated, in view of its application across the sciences and engineering.
For strings of symbols from $A$, we write $x_{1}^{n}$ for the block $x_{1}^{n}=(x_{1},x_{2},\ldots,x_{n})$, and similarly, for blocks of random variables, $X_{1}^{n}$ denotes $(X_{1},X_{2},\ldots,X_{n})$. A variable-rate code, or fixed-to-variable compressor, for strings of length $n$ from $A$ is an injective function $f_{n}:A^{n}\to\{0,1\}^{*}$, where $\{0,1\}^{*}$ denotes the set of all finite-length binary strings:
$\{0,1\}^{*}=\{\varnothing,0,1,00,01,10,11,000,001,\ldots\}.$
Writing $\ell(s)$ for the length of a string $s\in\{0,1\}^{*}$, the compressor $f_{n}$ maps strings $x_{1}^{n}\in A^{n}$ into binary strings $f_{n}(x_{1}^{n})$ of length $\ell(f_{n}(x_{1}^{n}))$ bits.
We are interested in the best achievable performance among all fixed-to-variable compressors. For a blocklength $n\geq 1$ and rate $R>0$, the best achievable excess-rate probability $\epsilon_{n}^{*}(R)$, is:
$\epsilon_{n}^{*}(R):=\min_{f_{n}}\mathbb{P}\big{(}\ell(f_{n}(X_{1}^{n}))\geq nR\big{)}.$
This minimum is achieved by an optimal compressor, $f_{n}^{*}$, which is independent of the rate $R$ [14]. Similarly, for a blocklength $n\geq 1$ and excess-rate probability $\epsilon>0$, the best achievable rate $R_{n}^{*}(\epsilon)$, is:
$R_{n}^{*}(\epsilon):=\inf\big{\{}R>0:\min_{f_{n}}\mathbb{P}\big{(}\ell(f_{n}(X_{1}^{n}))\geq nR\big{)}\leq\epsilon\big{\}}.$
Prefix-free codes. In the case of prefix-free codes, analogous fundamental limits, $\epsilon^{\sf p}_{n}(R)$ and $R_{n}^{\sf p}(\epsilon)$, can be defined, by restricting the corresponding minimisations to codes that satisfy the prefix condition. But $R_{n}^{\sf p}(\epsilon)$ is very tightly coupled with $R_{n}^{*}(\epsilon)$: As shown in [14, Theorem 1], we always have $R^{\sf p}_{n}(\epsilon)=R_{n}^{*}(\epsilon)+\frac{1}{n}$.
1.2 Background
When the excess-rate probability, $\epsilon$, is fixed, the problem of determining the first-order asymptotic behaviour of the optimal rate $R_{n}^{*}(\epsilon)$ essentially goes back to Shannon. We have,
$nR_{n}^{*}(\epsilon)=nH(P)+o(n)\;\;\mbox{bits},\quad\mbox{as}\;n\to\infty,$ (1)
where $H(P)=-\sum_{x}P(x)\log P(x)$ as usual denotes the entropy, and, throughout, ‘log’ denotes $\log_{2}$, the logarithm taken to base 2; see [14] for an extensive discussion and historical comments.
The first-order relation (1) was refined by Strassen [23] who claimed that, as $n\to\infty$,
$nR_{n}^{*}(\epsilon)=nH(P)+\sqrt{n}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2}\log n+O(1)\;\;\mbox{bits},$ (2)
[…]
When the excess-rate probability is not fixed but is instead required to decay to zero exponentially fast, the best rate that can be achieved turns out to be higher than the entropy. Specifically, for $\delta>0$ in an appropriate range, as $n\to\infty$,
$nR_{n}^{*}(2^{-n\delta})=nH(P_{\alpha^{*}})+o(n)\;\;\mbox{bits},$ (3)
for a specific $\alpha^{*}\in(0,1)$ depending on $\delta$, where, for each $\alpha\in(0,1)$, $P_{\alpha}$ denotes the tilted distribution,
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{\sum_{y\in A}P(y)^{\alpha}},\quad x\in A.$
The first-order asymptotic relation (3) is simply the “rate version” of the well-known error-exponents result that was established in the early works [9, 12] and in the present form by Blahut [4]. Indeed, $H(P_{\alpha^{*}})$ can be expressed as the inverse of the error-exponent function,
$\Delta(R):=\inf_{P^{\prime}:H(P^{\prime})\geq R}D(P^{\prime}\|P),$
via,
$H(P_{\alpha^{*}})=\Delta^{-1}(\delta)=\sup_{P^{\prime}:D(P^{\prime}\|P)\leq\delta}H(P^{\prime}),$
where $D(P^{\prime}\|P)$ is the relative entropy between $P^{\prime}$ and $P$.
1.3 Compression at pragmatic rates
The main contribution of the present work is a nonasymptotic refinement of the evaluation of the fundamental compression limit $R_{n}^{*}$, in the small excess-rate probability regime. In the same sense in which Strassen’s expansion (2) and the finite-$n$ bounds of [14] strengthen the classical Shannon asymptotic relation (1), we show it is possible to derive finite-$n$ bounds that provide a correspondingly finer and stronger version of the first-order asymptotic expansion (3) in the small excess-rate probability regime.
In Theorems 3.1 and 3.2 of Section 3 we provide explicit constants $C$ and $C^{\prime}$ such that,
$C^{\prime}\leq nR_{n}^{*}(2^{-n\delta})-\Big{[}nH(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\log n\Big{]}\leq C,$ (4)
for all $n$ greater than some explicit $N_{0}$. In view of this, we may call the expression,
[…]
the best pragmatic rate that can be achieved at blocklength $n$ with excess-rate probability no greater than $\epsilon=2^{-n\delta}$.
On account of the remark at the end of Section 1.1, the exact same results as in (4) hold in the case of codes with prefix constraints – that is, for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with the same constant $C^{\prime}$, and with $C+1$ in place of $C$.
Despite their technical nature, the approximations to $R_{n}^{*}$ and $R_{n}^{\sf p}$ provided by all these different approaches are of as much practical relevance as they are of mathematical interest. In particular, they each are useful in different regimes of the blocklength and the excess-rate probability requirements. For example, the Strassen-style approximation suggested by (2) is only relevant for moderate values of $\epsilon$. If $\epsilon$ is small, then the second term dominates and the approximation is no longer valid or useful.
[…]
Shannon: $R_{n}^{*}\approx H(P),$ (5)
Strassen: $R_{n}^{*}\approx H(P)+\frac{1}{\sqrt{n}}\sigma(P)Q^{-1}(\epsilon)-\frac{1}{2n}\log n,$ (6)
and in the ‘error exponents regime’,
Blahut: $R_{n}^{*}\approx H(P_{\alpha^{*}})$ (7)
Thms. 3.1-3.2: $R_{n}^{*}\approx{\cal R}_{n}=H(P_{\alpha^{*}})-\frac{1}{2n(1-\alpha^{*})}\log n.$ (8)
Table 1 shows representative results when the blocklength $n=50$, which clearly demonstrate that the pragmatic rate ${\cal R}_{n}(\epsilon)$ provides by far the most accurate estimate of the optimal rate $R_{n}^{*}(\epsilon)$ for the problem parameters considered.
$\epsilon$ $R_{n}^{*}(\epsilon)$ Shannon (5) Strassen (6) Blahut (7) ${\cal R}_{n}(\epsilon)$ (8)
0.00003 0.940 0.722 1.119 1.000 0.941
0.00010 0.940 0.722 1.086 0.997 0.936
0.00032 0.920 0.722 1.052 0.993 0.928
0.00093 0.900 0.722 1.017 0.987 0.917
0.00251 0.900 0.722 0.983 0.979 0.903
0.00626 0.880 0.722 0.948 0.969 0.888
0.01444 0.840 0.722 0.913 0.957 0.869
Table 1: Comparison between the true value of the optimal rate $R_{n}^{*}(\epsilon)$ and four different approximations. Clearly the approximation suggested by Theorems 3.1 and 3.2, given by the pragmatic rate ${\cal R}_{n}$, gives the best results in this regime.
1.4 Prior related work
Formally, the problem of lossless data compression is very closely related to binary hypothesis testing, especially in the case of fixed-rate compression. Indeed, there are close hypothesis-testing parallels to all the major data-compression steps outlined in Section 1.2, beginning with Stein’s lemma [5, 16, 6], its refinement by Strassen [23] and Polyanskiy-Poor-Verdú [22], and the error-exponent asymptotics of Hoeffding [11].
For hypothesis testing, sharper expansions in the error-exponents regime have been developed by numerous authors, including the work reported in [8, 1, 2, 25, 19, 20]. The results in this paper can be viewed as lossless compression analogues of the refined error-exponent bounds recently obtained for binary hypothesis testing in [18, 17].
For fixed-rate data compression, Csiszár and Longo [8] obtained an asymptotic expansion for the excess-rate probability, when the target rate is fixed. Our Theorems 3.1 and 3.2 can be viewed as finite-$n$ improvements of the “rate version” of these expansions, in the variable-rate case. Also, in a related but different direction, the best achievable rate in the “moderate deviations” regime was examined in [3].
Finally, we mention that extensive discussions of different aspects of the fundamental limits of lossless data compression can be found in Csiszár and Körner’s classic text, [7], Han’s book on information spectrum methods [10], and Tan’s monograph [24].
2 Preliminaries and auxiliary results
[…]
$P_{\alpha}(x)=\frac{P(x)^{\alpha}}{Z_{\alpha}},\quad x\in A,$ (9)
where the normalising constant $Z_{\alpha}=\sum_{x\in A}P(x)^{\alpha}$. The standard normal cumulative distribution function is denoted by $\Phi$, ‘$\log$’ denotes the logarithm taken to base 2, and the relative entropy between two p.m.f.s $P,Q$ on the same finite alphabet $A$ is $D(P\|Q)=\sum_{x\in A}P(x)\log[P(x)/Q(x)].$
2.1 Normal approximation, compression and hypothesis testing
Here we collect some classical normal approximation inequalities, a simple one-shot converse bound for data compression and some useful nonasymptotic results for binary hypothesis testing, all of which will be employed in the proofs of our main results in Section 3.
First, we recall the classical Berry-Esséen bound [15, 21] for sums of i.i.d. random variables.
Theorem 2.1 (Berry-Esséen bound)
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be i.i.d. random variables with mean $\mu=\mathbb{E}(Z_{1})$, variance $\sigma^{2}=\textup{Var}(Z_{1})$ and $\rho=\mathbb{E}(|Z_{1}-\mu|^{3})<\infty$. Then:
$\Bigg{|}\mathbb{P}\Bigg{(}\frac{\sqrt{n}}{\sigma}\Big{(}\frac{1}{n}\sum_{i=1}^{n}Z_{i}-\mu\Big{)}\leq x\Bigg{)}-\Phi(x)\Bigg{|}\leq\frac{\rho}{2\sigma^{3}\sqrt{n}},\hskip 14.45377ptx\in\mathbb{R},\hskip 5.05942ptn\geq 1.$
The following is a simple corollary of Theorem 2.1, cf. [22, 18, 17].
Lemma 2.2
Let $Z_{1},Z_{2},\ldots,Z_{n}$ be as in Theorem 2.1. Then,
$\mathbb{E}\Bigg{\{}\exp\Big{(}-\sum_{i=1}^{n}Z_{i}\Big{)}\mathbb{I}_{\{\sum_{i=1}^{n}Z_{i}\geq x\}}\Bigg{\}}\leq\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho}{\sigma^{2}}\Big{)}\frac{1}{\sigma\sqrt{n}}e^{-x},\hskip 14.45377ptx\in\mathbb{R},\hskip 5.05942ptn\geq 1.$
The following one-shot converse bound for data compression was established in [14].
Theorem 2.3 (One-shot converse)
Let $f^{*}$ be the optimal compressor for a finite-valued random variable $X$. Then, for any $k>0$ and $\tau>0$,
$\mathbb{P}[\ell(f^{*}(X))\geq k-1]>\mathbb{P}[-\log P(X)\geq k+\tau]-2^{-\tau}.$
The next two results were originally developed in the context of binary hypothesis testing. Lemma 2.4 [8] can be viewed as a “change of measure” result for hypothesis testing, and Theorem 2.5 [18, 17, 22] is a strong, finite-$n$ converse bound for hypothesis testing.
Lemma 2.4 (Change of measure)
Let $P,Q$ be two p.m.f.s with full support on $A$, and let $S_{n}$ be an arbitrary subset of $A^{n}$, for some $n\geq 1$. If,
$\frac{P^{n}(x_{1}^{n})}{Q^{n}(x_{1}^{n})}\geq\max_{y_{1}^{n}\in S_{n}}\frac{P^{n}(y_{1}^{n})}{Q^{n}(y_{1}^{n})}\quad\mbox{for all}\;x_{1}^{n}\not\in S_{n},$
then,
$P^{n}(S_{n})=\min_{S\subset A^{n}:Q^{n}(S^{c})\leq Q^{n}(S_{n}^{c})}P^{n}(S).$
Theorem 2.5 (Hypothesis testing converse)
Let $P\neq Q$ be p.m.f.s with full support on $A$, and for $n\geq 1$, $\epsilon\in(0,1)$, let,
$e_{1}^{*}(\epsilon)=\min_{S\subset A^{n}:Q^{n}(S^{c})\leq\epsilon}P^{n}(S).$
Then, for any $\Delta>0$ such that $\epsilon+\frac{B+\Delta}{\sqrt{n}}<1$,
$\log e_{1}^{*}(\epsilon)\geq-nD(Q\|P)-\sigma\sqrt{n}\Phi^{-1}\Big{(}\epsilon+\frac{B+\Delta}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n+\log\Delta,$
where $\sigma^{2}=\textup{Var}_{Q}(\log\frac{Q(X)}{P(X)})$, $\rho=\mathbb{E}_{Q}|\log\frac{Q(X)}{P(X)}-\mathbb{E}_{Q}(\log\frac{Q(X)}{P(X)})|^{3}$ and $B=\frac{\rho}{2\sigma^{3}}$.
2.2 Moments and derivatives identities
We close with some simple technical lemmas, stated without proof. Let $X\sim P$ where $P$ has full support on $A$, and recall the definition of $P_{\alpha}$ in (9). The proofs of Lemmas 2.62.9 follow by direct computation; see for example, [8] and [18, 17].
Lemma 2.6
Define the tilted second and third moments as:
$\sigma_{2,\alpha}^{2}$ $=\textup{Var}_{P_{\alpha}}\Big{(}\log_{e}\frac{P_{\alpha}(X)}{P(X)}\Big{)},$
$\rho_{2,\alpha}$ $=\mathbb{E}_{P_{\alpha}}\Big{|}\log_{e}\frac{P_{\alpha}(X)}{P(X)}-\mathbb{E}_{P_{\alpha}}\Big{(}\log_{e}\frac{P_{\alpha}(X)}{P(X)}\Big{)}\Big{|}^{3}\,,$
$\sigma_{3,\alpha}^{2}$ $=\textup{Var}_{P_{\alpha}}(\log_{e}P(X)),$
$\rho_{3,\alpha}$ $=\mathbb{E}_{P_{\alpha}}|\log_{e}P(X)-\mathbb{E}_{P_{\alpha}}(\log_{e}P(X))|^{3}.$
Then, for all $\alpha\in(0,1)$: $\sigma_{2,\alpha}^{2}=(1-\alpha)^{2}\sigma_{3,\alpha}^{2}$ and $\rho_{2,\alpha}=(1-\alpha)^{3}\rho_{3,\alpha}$.
Lemma 2.7
For all $\alpha\in(0,1)$ we have the following expressions for the derivatives of $D(P_{\alpha}\|P)$ and $\sigma^{2}_{3,\alpha}$:
$\frac{dD(P_{\alpha}\|P)}{d\alpha}$ $=(\alpha-1)\sigma_{3,\alpha}^{2}(\log e),$
$\frac{d^{2}D(P_{\alpha}\|P)}{d\alpha^{2}}$ $=(\log e)\sigma_{3,\alpha}^{2}+(\log e)(\alpha-1)\frac{d\sigma_{3,\alpha}^{2}}{d\alpha},$
$\frac{d\sigma_{3,\alpha}^{2}}{d\alpha}$ $=\mathbb{E}_{P_{\alpha}}\Big{[}\Big{(}\log_{e}P(X)-\mathbb{E}_{P_{\alpha}}[\log_{e}P(X)]\Big{)}^{3}\Big{]}.$
Lemma 2.8
$D(P_{\alpha}\|P)$ is a continuous, strictly decreasing function of $\alpha$, for $\alpha\in(0,1)$.
Lemma 2.9
For all $\alpha\in(0,1)$ we have the following expressions for the derivatives of $H(P_{\alpha})$:
$\frac{dH(P_{\alpha})}{d\alpha}$ $=-(\log e)\alpha\sigma_{3,\alpha}^{2},$
$\frac{d^{2}H(P_{\alpha})}{d\alpha^{2}}$ $=-(\log e)\Big{[}\sigma_{3,\alpha}^{2}+\alpha\frac{d\sigma_{3,\alpha}^{2}}{d\alpha}\Big{]}.$
3 Main results
We will now proceed with the main results of this paper, namely the development of non-asymptotic achievability and converse bounds on the optimal compression rate, $R_{n}^{*}(2^{-n\delta})$, when the excess-rate probability decays exponentially fast.
3.1 Achievability
Theorem 3.1 (Achievability)
Consider a memoryless source $X$ with p.m.f. $P$ with full support on $A$. Let $\delta\in(0,D(U\|P))$, where $U$ is the uniform distribution over $A$. Then, for any $n\geq 1$,
$R_{n}^{*}(2^{-n\delta})\leq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}+\frac{c}{n},$ (10)
where $P_{\alpha}$ is defined in (9), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and,
$c=\log\Bigg{(}\frac{1}{\sigma_{1}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1}}{\sigma_{1}^{2}}\Big{)}\Bigg{)}+\frac{\alpha^{*}}{1-\alpha^{*}}\log\Bigg{(}\frac{1}{\sigma_{2}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2}}{\sigma_{2}^{2}}\Big{)}\Bigg{)},$
where $\sigma_{1}^{2}=\textup{Var}_{P_{\alpha^{*}}}(\log_{e}P_{\alpha^{*}}(X))$, $\rho_{1}=\mathbb{E}_{P_{\alpha^{*}}}|\log_{e}P_{\alpha^{*}}(X)-\mathbb{E}_{P_{\alpha^{*}}}(\log_{e}P_{\alpha^{*}}(X))|^{3}$, $\sigma_{2}^{2}=\sigma_{2,\alpha^{*}}^{2}$ and $\rho_{2}=\rho_{2,\alpha^{*}}$, with $\sigma_{2,\alpha}^{2}$ and $\rho_{2,\alpha}$ defined as in Lemma 2.6.
Remark. As noted in the introduction, the exact bound (10) remains valid in the case of codes with prefix constraints – i.e., for $R_{n}^{\sf p}$ in place of $R_{n}^{*}$ – with $c+1$ in place of $c$.
Proof. The proof is based on an explicit construction of a compressor $C_{n}$. To that end, let $10<\beta_{n}<$ 1 be a constant that will be chosen later, and define the set,
$E_{n}=\{x_{1}^{n}\in A^{n}:\log P^{n}(x_{1}^{n})\geq\log\beta_{n}\}.$ (11)
Write $M_{X}(\beta_{n})=|E_{n}|$ for the cardinality of $E_{n}$, and suppose that,
$P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$ (12)
To obtain the desired compressor on $A^{n}$, first we order the strings $x_{1}^{n}\in E_{n}$ in order of decreasing probability $P^{n}(x_{1}^{n})$, breaking ties arbitrarily, and define $C_{n}$ via:
$\ell(C_{n}(x_{1}^{n}))=\begin{cases}\lfloor\log k\rfloor&\text{if $x_{1}^{n}$ is the $k^{th}$ element of $E_{n}$},\\ \lceil\log|A|^{n}\rceil&\text{if $x_{1}^{n}\not\in E_{n}$}.\end{cases}$
Note that the longest codeword $C_{n}(x_{1}^{n})$ among all strings in $E_{n}$ has length $\lfloor\log M_{X}(\beta_{n})\rfloor$. Therefore, for any fixed $\xi\in(0,1)$,
$\mathbb{P}[\ell(C_{n}(X_{1}^{n}))\geq\lfloor\log M_{X}(\beta_{n})\rfloor+\xi]\leq P^{n}(E_{n}^{c})\leq 2^{-n\delta}.$
Hence, by definition, $nR_{n}^{*}(2^{-n\delta})\leq\log M_{X}(\beta_{n})+\xi$, and since $\xi\in(0,1)$ was arbitrary,
$R_{n}^{*}(2^{-n\delta})\leq\frac{\log M_{X}(\beta_{n})}{n}.$
The remainder of the proof is devoted to showing that there is an appropriate $\beta_{n}$ such that (12) holds, and evaluating the rate $(1/n)\log M_{X}(\beta_{n})$ for this choice.
Let $\alpha\in(0,1)$ arbitrary. We begin by observing that we can express,
$M_{X}(\beta_{n})$ $=\sum_{x_{1}^{n}\in E_{n}}\frac{P^{n}_{\alpha}(x_{1}^{n})}{P^{n}_{\alpha}(x_{1}^{n})}$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp(-\log_{e}P_{\alpha}^{n}(x_{1}^{n}))\mathbb{I}_{E_{n}}(x_{1}^{n})$
$=\mathbb{E}_{P_{\alpha}^{n}}\Big{[}\exp\Big{(}-\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\Big{)}\mathbb{I}\Big{\{}\sum_{i=1}^{n}\log_{e}P(X_{i})\geq\log_{e}\beta_{n}\Big{\}}\Big{]}.$
We can now apply the version of the Berry-Esséen bound in Lemma 2.2, with $Z_{i}=\log_{e}P_{\alpha}(X_{i})$ and $x=\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}$, to obtain that:
$M_{X}(\beta_{n})$ $=\mathbb{E}_{P_{\alpha}^{n}}\Big{[}\exp\Big{(}-\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\Big{)}\mathbb{I}\Big{\{}\sum_{i=1}^{n}\log_{e}P_{\alpha}(X_{i})\geq\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}\Big{\}}\Big{]}$
$\leq\frac{1}{\sigma_{1,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big{)}\frac{1}{\sqrt{n}}\exp(-(\alpha\log_{e}\beta_{n}-n\log_{e}Z_{\alpha}))$
$=\frac{1}{\sigma_{1,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big{)}\frac{Z_{\alpha}^{n}}{\beta_{n}^{\alpha}\sqrt{n}},$ (13)
where $\sigma_{1,\alpha}^{2}=\text{Var}_{P_{\alpha}}(\log_{e}P_{\alpha}(X))$ and $\rho_{1,\alpha}=\mathbb{E}_{P_{\alpha}}|\log_{e}P_{\alpha}(X)-\mathbb{E}_{P_{\alpha}}(\log_{e}P_{\alpha}(X))|^{3}$.
In order to select appropriate values for $\beta_{n}$ and $\alpha$, we examine $P^{n}(E_{n}^{c})$:
$P^{n}(E_{n}^{c})$ $=\sum_{x_{1}^{n}\in A^{n}}P^{n}(x_{1}^{n})\mathbb{I}_{E_{n}^{c}}(x_{1}^{n})$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}\log_{e}\frac{P^{n}(x_{1}^{n})}{P_{\alpha}^{n}(x_{1}^{n})}\Bigg{)}\mathbb{I}\{\log_{e}P^{n}(x_{1}^{n})<\log_{e}\beta_{n}\}$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}\mathbb{I}\Bigg{\{}\sum_{i=1}^{n}\log_{e}P(x_{i})<\log_{e}\beta_{n}\Bigg{\}}.$
Rearranging, we can further express,
$P^{n}(E_{n}^{c})$ $=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}$
$\hskip 72.26999pt\mathbb{I}\Bigg{\{}n\log_{e}Z_{\alpha}+\sum_{i=1}^{n}\log_{e}P^{1-\alpha}(x_{i})<n\log_{e}Z_{\alpha}+(1-\alpha)\log_{e}\beta_{n}\Bigg{\}}$
$=\sum_{x_{1}^{n}\in A^{n}}P_{\alpha}^{n}(x_{1}^{n})\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}\Bigg{)}$
$\hskip 72.26999pt\mathbb{I}\Bigg{\{}\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(x_{i})}{P(x_{i})}>-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}\Bigg{\}},$
so that,
$P^{n}(E_{n}^{c})$ $=\mathbb{E}_{P_{\alpha}^{n}}\Bigg{\{}\exp\Bigg{(}-\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(X_{i})}{P(X_{i})}\Bigg{)}$
$\hskip 72.26999pt\mathbb{I}\Bigg{\{}\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha}(X_{i})}{P(X_{i})}>-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}\Bigg{\}}\Bigg{\}}.$
Applying Lemma 2.2 again, with $Z_{i}=\log_{e}[P_{\alpha}(X_{i})/P(X_{i})]$ and $x=-n\log_{e}Z_{\alpha}-(1-\alpha)\log_{e}\beta_{n}$, yields,
$P^{n}(E_{n}^{c})$ $\leq\frac{1}{\sigma_{2,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big{)}\frac{1}{\sqrt{n}}\exp\Big{(}n\log_{e}Z_{\alpha}+(1-\alpha)\log_{e}\beta_{n}\Big{)}$
$=\frac{1}{\sigma_{2,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big{)}\frac{Z_{\alpha}^{n}\beta_{n}^{1-\alpha}}{\sqrt{n}}.$ (14)
Now, setting,
$\log\beta_{n}=\frac{1}{1-\alpha}\Bigg{(}-n\delta+\log\Bigg{(}\frac{1}{\frac{1}{\sigma_{2,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big{)}\frac{Z_{\alpha}^{n}}{\sqrt{n}}}\Bigg{)}\Bigg{)},$
immediately yields (12). Note that, $\beta_{n}>0$ by (14), while (11) and (12) imply $\beta_{n}<1$.
As a result, the upper bound for $M_{X}(\beta_{n})$ in (13) becomes:
$\log M_{X}(\beta_{n})$ $\leq\log\Bigg{(}\frac{1}{\sigma_{1,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big{)}\frac{Z_{\alpha}^{n}}{\beta_{n}^{\alpha}\sqrt{n}}\Bigg{)}$
$=\log\Bigg{(}\frac{1}{\sigma_{1,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big{)}\Bigg{)}+n\log Z_{\alpha}-\frac{1}{2}\log n$
$\hskip 36.135pt+\frac{\alpha n\delta}{1-\alpha}-\frac{\alpha}{1-\alpha}\log\Bigg{(}\frac{1}{\frac{1}{\sigma_{2,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big{)}\frac{Z_{\alpha}^{n}}{\sqrt{n}}}\Bigg{)}$
$=\frac{n}{1-\alpha}(\alpha\delta+\log Z_{\alpha})-\frac{1}{2(1-\alpha)}\log n$
$\hskip 36.135pt+\log\Bigg{(}\frac{1}{\sigma_{1,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{1,\alpha}}{\sigma_{1,\alpha}^{2}}\Big{)}\Bigg{)}+\frac{\alpha}{1-\alpha}\log\Bigg{(}\frac{1}{\sigma_{2,\alpha}}\Big{(}\frac{1}{\sqrt{2\pi}}+\frac{\rho_{2,\alpha}}{\sigma_{2,\alpha}^{2}}\Big{)}\Bigg{)}.$ (15)
We finally need to select $\alpha$. Since $D(P_{\alpha}\|P)$ is continuous and strictly decreasing in $\alpha$ (Lemma 2.8), and since $\delta\in(D(P_{1}\|P)),D(P_{0}\|P))=(0,D(U\|P))$, there exists a unique $\alpha^{*}\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$. For this value of $\alpha$,
$\alpha^{*}\delta+\log Z_{\alpha^{*}}$ $=\alpha^{*}\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)}{P(x)}+\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)Z_{\alpha^{*}}}{P_{\alpha^{*}}(x)}$
$=\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P_{\alpha^{*}}(x)^{\alpha^{*}}}{P(x)^{\alpha^{*}}}+\sum_{x\in A}P_{\alpha^{*}}(x)\log\frac{P(x)^{\alpha^{*}}}{P_{\alpha^{*}}(x)}$
$=\sum_{x\in A}P_{\alpha^{*}}(x)\log\Big{(}P_{\alpha^{*}}(x)^{(\alpha^{*}-1)}\Big{)}$
$=(1-\alpha^{*})H(P_{\alpha^{*}}).$
Substituting this in (15) yields (10), as required.     $\Box$
3.2 Converse
Theorem 3.2 (Converse)
Consider a memoryless source $X$ with p.m.f. $P$ with full support on $A$. Let $\delta\in(0,D(U\|P))$, where $U$ is the uniform distribution over $A$. Then for any $n>N_{0}$,
$R_{n}^{*}(2^{-n\delta})\geq H(P_{\alpha^{*}})-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{C}{n},$ (16)
where $P_{\alpha}$ is as defined in (9), $\alpha^{*}$ is the unique $\alpha\in(0,1)$ for which $\delta=D(P_{\alpha^{*}}\|P)$, and,
$N_{0}$ $=\max\Bigg{\{}4.4\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}^{2},\frac{4(1+q+r)^{2}}{p^{2}},\frac{2(1+q+r)}{p(1-\alpha^{*})},N_{1},N_{2}\Bigg{\}},$
$C$ $=\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})+1,$
where $\sigma_{3,\alpha}$ and $\rho_{3,\alpha}$ are as defined in Lemma 2.6, and the remaining constants are given by:
$\sigma_{3}^{2}=\sigma_{3,\alpha^{*}}^{2},\;\;\tilde{\sigma_{3}}^{2}=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2},\;\;\hat{\rho_{3}}=\sup_{\alpha\in(0,1)}\rho_{3,\alpha},$
$p=(\log e)(1-\alpha^{*})\sigma_{3}^{2},\;\;q=\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}),\;\;r=19(\log e)(1-\alpha^{*})\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}},$
$N_{1}=\min\big{\{}n\geq 8\;:\;\log n\leq p\sqrt{n}\big{\}},\;\;N_{2}=\min\big{\{}n\geq 3\;:\;\log n\leq p(1-\alpha^{*})n\big{\}}.$
Remark. Again, in view of the remark at the end of Section 1.1, the bound (16) remains valid for prefix free codes exactly as stated.
Proof. Taking $k=nR$, for some $R>0$, letting $\tau=n\delta$, and replacing $X$ with $X_{1}^{n}$ (and $f^{*}$ with $f_{n}^{*}$) in the one-shot converse in Theorem 2.3, yields:
$\mathbb{P}\Big{[}\ell(f_{n}^{*}(X_{1}^{n}))\geq n\Big{(}R-\frac{1}{n}\Big{)}\Big{]}>\mathbb{P}[-\log P^{n}(X_{1}^{n})\geq n(R+\delta)]-2^{-n\delta}.$
Therefore, if we can find a rate $R$ for which,
$\mathbb{P}[\log P^{n}(X_{1}^{n})\leq-n(R+\delta)]\geq 2\times 2^{-n\delta},$
this would imply that $\mathbb{P}[\ell(f_{n}^{*}(X_{1}^{n}))\geq n(R-1/n)]>2^{-n\delta}$, which would in turn give,
$R_{n}^{*}(2^{-n\delta})\geq R-\frac{1}{n}.$ (17)
Writing,
$S_{n}=\{x_{1}^{n}\in A^{n}:\log P^{n}(x_{1}^{n})\leq-n(R+\delta)\},$
our objective then is to find an appropriate rate $R$ for which,
$P^{n}(S_{n})\geq 2\times 2^{-n\delta}.$ (18)
Much of the remainder of the proof will be devoted to establishing (18).
First we will use Lemma 2.4 to express $P^{n}(S_{n})$ in terms of $P_{\alpha_{n}}^{n}(S_{n}^{c})$ for some $\alpha_{n}\in(0,1)$ that will be chosen later. Indeed, since,
$S_{n}$ =$=\Bigg{\{}x_{1}^{n}\in A^{n}:\sum_{i=1}^{n}\log P^{1-\alpha_{n}}(x_{i})\leq-(1-\alpha_{n})n(R+\delta)\Bigg{\}}$
=$=\Bigg{\{}x_{1}^{n}\in A^{n}:\log\frac{P^{n}(x_{1}^{n})}{P^{n}_{\alpha_{n}}(x_{1}^{n})}\leq-(1-\alpha_{n})n(R+\delta)+n\log Z_{\alpha_{n}}\Bigg{\}},$ (19)
we have that,
$\frac{P^{n}(x_{1}^{n})}{P_{\alpha_{n}}^{n}(x_{1}^{n})}\geq\max_{x_{1}^{n}\in S_{n}}\frac{P^{n}(x_{1}^{n})}{P_{\alpha_{n}}^{n}(x_{1}^{n})}\quad\mbox{for all}\;x_{1}^{n}\not\in S_{n},$
so we can apply Lemma 2.4 to obtain that,
$P^{n}(S_{n})=\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq P_{\alpha_{n}}^{n}(S_{n}^{c})}P^{n}(S).$ (20)
The next step will be an application of the Berry-Esséen bound to obtain an upper bound on $P_{\alpha_{n}}^{n}(S_{n}^{c})$. From (19) we get,
$P_{\alpha_{n}}^{n}(S_{n}^{c})=P_{\alpha_{n}}^{n}\Bigg{(}\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha_{n}}(x_{i})}{P(x_{i})}<\frac{(1-\alpha_{n})n(R+\delta)-n\log Z_{\alpha_{n}}}{\log e}\Bigg{)}.$ (21)
Choosing $R$ as,
$R=\frac{1}{1-\alpha_{n}}\Big{[}\mathbb{E}_{P_{\alpha_{n}}}\Big{(}\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big{)}+\log Z_{\alpha_{n}}\Big{]}-\delta,$ (22)
we have,
$\frac{(1-\alpha_{n})n(R+\delta)-n\log Z_{\alpha_{n}}}{\log e}=n\mathbb{E}_{P_{\alpha_{n}}}\Big{(}\log_{e}\frac{P_{\alpha_{n}}(X)}{P(X)}\Big{)},$
and the Berry-Esséen in Theorem 2.1 applied to (21) yields,
$P_{\alpha_{n}}^{n}(S_{n}^{c})=P_{\alpha_{n}}^{n}\Bigg{(}\sum_{i=1}^{n}\log_{e}\frac{P_{\alpha_{n}}(x_{i})}{P(x_{i})}<n\mathbb{E}_{P_{\alpha_{n}}}\Big{(}\log_{e}\frac{P_{\alpha_{n}}(X)}{P(X)}\Big{)}\Bigg{)}\leq\Phi(0)+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}.$
Using Lemma 2.6, the bound becomes,
$P_{\alpha_{n}}^{n}(S_{n}^{c})\leq\frac{1}{2}+\frac{\frac{\rho_{2,\alpha_{n}}}{(1-\alpha_{n})^{3}}}{\frac{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}{(1-\alpha_{n})^{3}}}=\frac{1}{2}+\frac{\rho_{3,\alpha_{n}}}{2\sigma_{3,\alpha_{n}}^{3}\sqrt{n}}\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}},$ (23)
where $\sigma_{3,\alpha_{n}}$ and $\rho_{3,\alpha_{n}}$ are as in Lemma 2.6, $\tilde{\sigma_{3}}^{2}:=\inf_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$ and $\hat{\rho_{3}}:=\sup_{\alpha\in(0,1)}\rho_{3,\alpha}$.
Using this bound, from (20) we get,
$P^{n}(S_{n})\geq\min_{S\subset A^{n}:P_{\alpha_{n}}^{n}(S^{c})\leq\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}}P^{n}(S).$
Then, using the hypothesis testing converse in Theorem 2.5, with $P_{\alpha_{n}}$ in the place of $Q$, $\epsilon=\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}$, and $\Delta=1$, yields,
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (24)
$\geq-nD(P_{\alpha_{n}}\|P)-\sigma_{2,\alpha_{n}}(\log e)\sqrt{n}\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}-\frac{1}{2}\log n,$ (25)
as long as $\frac{1}{2}+\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}<1$, i.e., for $n>4(\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}}+\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}+1)^{2}$. Note that $\frac{\hat{\rho_{3}}}{2\tilde{\sigma_{3}}^{3}}\geq\frac{\rho_{2,\alpha_{n}}}{2\sigma_{2,\alpha_{n}}^{3}}$ by (23), which justifies (25) and also means that (24) still holds when $n>4(\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1)^{2}$.
Recall that the existence of a unique $\alpha^{*}$ as in the statement of the theorem is already established in Theorem 3.1. In order to bound $\log P^{n}(S_{n})$ below further, we will use simple Taylor expansions to get bounds on $D(P_{\alpha_{n}}\|P)$, $\sigma_{2,\alpha_{n}}$ and the $\Phi^{-1}$ term in (25). First we expand $D(P_{\alpha_{n}}\|P)$ around $\alpha^{*}$. Using the expressions for the derivatives of $D(P_{\alpha}\|P)$ in Lemma 2.7, gives,
$D(P_{\alpha_{n}}\|P)$ $=D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)$
$\quad+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Bigg{(}\sigma_{3,u}^{2}+(u-1)\mathbb{E}_{P_{u}}\Big{[}\Big{(}\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big{)}^{3}\Big{]}\Bigg{)},$ (26)
for some $u\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha^{*}\leq\alpha_{n}$, and where $\sigma_{3}=\sigma_{3,\alpha^{*}}$. Letting $\hat{\sigma}_{3}^{2}=\sup_{\alpha\in(0,1)}\sigma_{3,\alpha}^{2}$, we have $\sigma_{3,u}^{2}\leq\hat{\sigma}_{3}^{2}$, and we observe that $|u-1|\leq 1$ and that,
$\Big{|}\mathbb{E}_{P_{u}}\Big{[}\Big{(}\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big{)}^{3}\Big{]}\Big{|}\leq\mathbb{E}_{P_{u}}\Big{|}\log_{e}P(X)-\mathbb{E}_{P_{u}}(\log_{e}P(X))\Big{|}^{3}\leq\hat{\rho}_{3}.$
Therefore,
$D(P_{\alpha_{n}}\|P)\leq D(P_{\alpha^{*}}\|P)+(\alpha_{n}-\alpha^{*})(\alpha^{*}-1)\sigma_{3}^{2}(\log e)+\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3}).$ (27)
Next, since $\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}$ by Lemma 2.6, we can expand $\sigma_{3,\alpha_{n}}$ instead of $\sigma_{2,\alpha_{n}}$ around $\alpha^{*}$, yielding,
$\sigma_{3,\alpha_{n}}^{2}=\sigma_{3}^{2}+(\alpha_{n}-\alpha^{*})\mathbb{E}_{P_{w}}\Big{[}\Big{(}\log_{e}P(X)-\mathbb{E}_{P_{w}}(\log_{e}P(X))\Big{)}^{3}\Big{]}\leq\sigma_{3}^{2}+\hat{\rho}_{3},$
for some $w\in[\alpha^{*},\alpha_{n}]$, assuming $\alpha_{n}-\alpha^{*}\leq 1$. Hence,
$\sigma_{2,\alpha_{n}}=(1-\alpha_{n})\sigma_{3,\alpha_{n}}\leq(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}.$ (28)
And for the third term, we expand $\Phi^{-1}$ around $1/2$ to obtain that,
$\Phi^{-1}\Big{(}\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}$ $=\Phi^{-1}\Big{(}\frac{1}{2}\Big{)}+\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}\Big{)}\frac{1}{\phi(\Phi^{-1}(x))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(\Phi^{-1}(\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}}))}$
$\leq\frac{1}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}\frac{1}{\phi(2)}\leq\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)},$ (29)
for some $x\in(\frac{1}{2},\frac{1}{2}+\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}\sqrt{n}}+\frac{1}{\sqrt{n}})$, where (29) holds as long as:
$n\geq 4.4\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}^{2}>\frac{(\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1)^{2}}{(\Phi(2)-1/2)^{2}}.$
Finally, substituting the bounds (27), (28) and (29) into (25), gives,
$\log P^{n}(S_{n})$ $\geq-nD(P_{\alpha^{*}}\|P)-n(\alpha_{n}-\alpha^{*})(\log e)(\alpha^{*}-1)\sigma_{3}^{2}-\frac{n\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+\hat{\rho}_{3})$
$\hskip 43.36243pt-(1-\alpha^{*})\sqrt{\sigma_{3}^{2}+\hat{\rho}_{3}}\times(\log e)\sqrt{n}\times\frac{19}{\sqrt{n}}\Big{(}\frac{\hat{\rho_{3}}}{\tilde{\sigma_{3}}^{3}}+1\Big{)}-\frac{1}{2}\log n$
$\geq-n\delta+pn(\alpha_{n}-\alpha^{*})-qn(\alpha_{n}-\alpha^{*})^{2}-r-\frac{1}{2}\log n,$ (30)
where $\delta,p,q$ and $r$ are as in the statement of the theorem.
Now we are in a position to choose $\alpha_{n}$. Letting,
$\alpha_{n}=\alpha^{*}+\frac{1}{2p}\frac{\log n}{n}+\frac{1+q+r}{p}\frac{1}{n},$ (31)
simple algebra shows that we have,
$pn(\alpha_{n}-\alpha^{*})-qn(\alpha_{n}-\alpha^{*})^{2}-r-\frac{1}{2}\log n\geq 1,$
as long as $n$ is large enough so that,
$\frac{\log n}{\sqrt{n}}\leq p\quad\mbox{and}\quad n\geq\frac{4(1+q+r)^{2}}{p^{2}},$ (32)
in which case (30) becomes $\log P^{n}(S_{n})\geq-n\delta+1$, which is exactly our desired bound (18). To ensure its validity, we must also ensure that the earlier conditions we applied – namely, that $\alpha_{n}\geq\alpha^{*}$, that $\alpha_{n}-\alpha^{*}\leq 1$ and that $\alpha_{n}\in(0,1)$ – are also satisfied. The inequalities $\alpha_{n}>0$ and $\alpha_{n}\geq\alpha^{*}$ follow from (31), $\alpha_{n}-\alpha^{*}<1$ follows from (31) and (32), and finally, if we further take $n$ large enough that
$\frac{\log n}{n}<p(1-\alpha^{*})\quad\mbox{and}\quad n>\frac{2(1+q+r)}{p(1-\alpha^{*})},$
then $\alpha_{n}<1$ easily follows from (31). To summarize, taking $n>N_{0}$ as in the statement of the theorem ensures the validity of our key bound (18).
The final step of the proof consists of showing that, with $R$ chosen as in (22) and with $\alpha_{n}$ as in (31), the rate $R$ can be bounded below so that, in combination with (17), we will obtain the desired bound (16) in the statement of the theorem.
To that end, starting from the definition of $R$ in (22), we can re-write,
$R$ $=\mathbb{E}_{P_{\alpha_{n}}}\Big{(}\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big{)}+\frac{1}{1-\alpha_{n}}\Big{[}\alpha_{n}\mathbb{E}_{P_{\alpha_{n}}}\Big{(}\log\frac{P_{\alpha_{n}}(X)}{P(X)}\Big{)}+\log Z_{\alpha_{n}}\Big{]}-\delta$
$=D(P_{\alpha_{n}}\|P)+\frac{1}{1-\alpha_{n}}\Big{[}\alpha_{n}D(P_{\alpha_{n}}\|P)+(1-\alpha_{n})H(P_{\alpha_{n}})-\alpha_{n}D(P_{\alpha_{n}}\|P)\Big{]}-\delta$
$=D(P_{\alpha_{n}}\|P)+H(P_{\alpha_{n}})-\delta.$ (33)
To bound this below, we will expand $D(P_{\alpha_{n}}\|P)$ and $H(P_{\alpha_{n}})$ around $\alpha^{*}$. For $D(P_{\alpha_{n}}\|P)$, continuing from the Taylor expansion in (26) we have,
$D(P_{\alpha_{n}}\|P)$ $\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (34)
$\geq\delta-(\alpha_{n}-\alpha^{*})(1-\alpha^{*})\sigma_{3}^{2}(\log e)-\frac{\log e}{2n}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}$ (35)
$=\delta-\frac{\log n}{2n}-\frac{1+q+r}{n}-\frac{\log e}{2n}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|},$ (36)
where (34) holds because the right-hand side is a lower bound to (26) since $\sigma_{3,u}^{2}\geq\tilde{\sigma_{3}}^{2}$, $1-u\leq 1-\alpha^{*}$ and by the definition of $\hat{\rho_{3}}$; (35) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (36) follows by substituting the expression of $\alpha_{n}$ in (31). Similarly expanding $H(P_{\alpha_{n}})$ around $\alpha^{*}$ and using the expressions for the derivatives of $H(P_{\alpha})$ in Lemma 2.9,
$H(P_{\alpha_{n}})$ $=H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e$
$\hskip 36.135pt-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}\Bigg{(}\sigma_{3,v}^{2}+v\mathbb{E}_{P_{v}}\Big{[}\Big{(}\log_{e}P(X)-\mathbb{E}_{P_{v}}(\log_{e}P(X))\Big{)}^{3}\Big{]}\Bigg{)}$ (37)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2}(\alpha_{n}-\alpha^{*})^{2}(\hat{\sigma}_{3}^{2}+1\times\hat{\rho_{3}})$ (38)
$\geq H(P_{\alpha^{*}})-(\alpha_{n}-\alpha^{*})\alpha^{*}\sigma_{3}^{2}\log e-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})$ (39)
$=H(P_{\alpha^{*}})-\frac{\alpha^{*}}{2(1-\alpha^{*})}\frac{\log n}{n}-\frac{\alpha^{*}(1+q+r)}{1-\alpha^{*}}\frac{1}{n}-\frac{\log e}{2n}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}}),$ (40)
where (37) holds for some $v\in[\alpha^{*},\alpha_{n}]$; (38) holds since $\sigma_{3,u}^{2}\leq\hat{\sigma_{3}}^{2}$, $v\leq 1$ and by the definition of $\hat{\rho_{3}}$; (39) follows from the definition of $\alpha_{n}$ and the fact that $n$ is taken sufficiently large as above; and (40) follows by substituting the expression of $\alpha_{n}$ in (31).
Substituting the bounds (36) and (40) into (33), gives,
$R\geq H(P_{\alpha^{*}})$ $-\frac{1}{2(1-\alpha^{*})}\frac{\log n}{n}$
$-\Bigg{\{}\frac{1+q+r}{1-\alpha^{*}}+\frac{\log e}{2}\Big{|}\tilde{\sigma_{3}}^{2}-(1-\alpha^{*})\hat{\rho_{3}}\Big{|}+\frac{\log e}{2}(\hat{\sigma}_{3}^{2}+\hat{\rho_{3}})\Bigg{\}}\times\frac{1}{n},$
and combining this with (17) yields (16) and completes the proof.     $\Box$